OpenStudy (anonymous):

Choose the correct description of the graph of the inequality -6x 30. Open circle on 5, shading to the left. Closed circle on 5, shading to the left. Open circle on -5, shading to the right. Closed circle on -5, shading to the right.

OpenStudy (anonymous):

someone please help

OpenStudy (debbieg):

You'll need to share with us what the inequality symbol is. -6x ??? 30 But basically, solve just like you would an Equality, with one exception: when you MULTIPLY or DIVIDE by a NEGATIVE number (as you have to do here) you have to flip the direction of the inequality. The only other part you need to know is that, if it is a STRICT inequality (> or <), then the solution set will NOT INCLUDE the endpoint - hence, the graph gets an OPEN CIRCLE at the endpoint. If it is a WEAK inequality (\(\ge\), \(\le\)), then the solution set DOES include the endpoint - hence, the graph gets a CLOSED circle.

OpenStudy (debbieg):

Eg, if I have: \(x \ge3\) , 3 IS in that solution set (since it makes the inequality true). So I have a closed circle on 3, and extend to all x's greater than 3, so to the right/ If I have: \(x >3\) it is almost the same, except with an open circle on 3 (since 3 does NOT make the inequality true). For \(x \le3\) or \(x < 3\), same as above, except the solution set includes all x's LESS THAN 3 so the graph extends to the left.

OpenStudy (anonymous):

≤ thats the inequality symbol

OpenStudy (anonymous):

I thought it would the last one closed circle on -5, shading to the right.

OpenStudy (debbieg):

\(\Large -6x \le 30\) So divide by -6 (and hence flip the inequality): \(\Large x \ge -5\) So yes, you're exactly right. :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (debbieg):

you're welcome. :)